3) Determine the values below using the graphs of fand g shown. If a value is not defined, state not defined. Assume the ends of fand g continue the shape you see beyond the graph. Ax) c) 2-5(g(0)-3) (Read carefully.) d) 9(3) a e) g-2) 6-2 9) R-3)+9(5) h) For what values of x does g(x)-2? i) For what values of x is fdecreasing? D For what values of x is g increasing? k) Give three values that g(x) can never be? (Is this asking for an x-value or a y-value?) l) Describe all the possible output values of fin a sentence. 4) What does the notation fx) really represent? Is it a function? An x-value? A y-value? An expression? (for what?) A complete equation? (for what)? Explain.
Answer #1

3. (a) f(0) = 0

(b). g(0) = -2

(c ).2 -5(g(0)-3) = 2-5(-2-3) = 2-5(-5) = 2+25 = 27.

(d). g(3) is undefined as there is a hole in the graph of g(cx) at = 3.

(e). g(-2) = -2.

(f). f(1)+g(1) = 5+(-2) = 3.

(g). f(-3)+g(5) = ? f(-3) = -3. The graph of g(x) is truncated beyond x = 4.

(h). g(x) = -2 for x = -2.

(i). f is decreasing when x is in the interval [-5,-2)